# Direct $N$-body simulations of rotating and extremely massive Population III star clusters

### 31 Mar 2026 v1

@@ 2 Methods
We use Nbody6++GPU (Spurzem, 1999; Nitadori & Aarseth, 2012; Wang et al., 2015) for the dynamical evolution of the Pop III clusters. Rotating initial conditions follow Kamlah et al. (2022b) and are generated with McLuster (Kuepper et al., 2011; Leveque et al., 2022b) and Fopax (Einsel & Spurzem, 1999); details are given in Appendix C.2.
Pop III stellar evolution is treated with the low-metallicity recipes of Tanikawa et al. (2020) at $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, without radiation-, pulsation- (e.g. Nakauchi et al., 2020), or rotation-driven (e.g. Liu et al., 2021b) wind mass loss; further details are in Appendix C.3. Fitting formulae for relativistic merger recoil kick velocity and remnant mass and spin follow Campanelli et al. (2007); Lousto et al. (2012); Jiménez-Forteza et al. (2017), as summarized in Appendix C.4. We refer to BHs with masses above 100 $\mathrm{M}_{\odot}$ as IMBHs.
@@ 3 Initial conditions
The mass function of Pop III star clusters is highly uncertain. Formation of Pop III star clusters has been simulated in cosmological mini-halos (c.f., e.g., Stacy et al., 2016), and corresponding star-cluster simulations have covered masses from $10^{3}$ to $10^{5}$ $\mathrm{M}_{\odot}$ (Liu et al., 2025; Wu et al., 2025). However, Kimm et al. (2016) propose the formation of more massive Pop III clusters without dark matter directly from cooling gas clouds, with masses up to $5\cdot 10^{5}\mathrm{M}_{\odot}{}$, and Liu et al. (2021a) find cluster masses reaching this value. Stacy et al. (2016) also discuss a possible Pop III cluster with mass as high as $10^{7}\mathrm{M}_{\odot}{}$. In our simulations we explore the very high-mass end (over $10^{5}$ objects, with 1% binaries and a top-heavy IMF; see Table 1) for two reasons: first, to extend previously published simulations to higher masses; and second, to probe very high initial densities that may promote rapid formation of massive stars and massive black holes (Vergara et al., 2025b).
@@
We start from an unsegregated, unfractal King model (King, 1966; Goodwin & Whitworth, 2004) with $W_{0}=6.0$, initially in virial equilibrium. The half-mass radius is set to $r_{\mathrm{h}}$=1 pc. As outlined in Appendix C.2, the initial model from McLuster is then redistributed with a rotating King model, which are more compact than their non-rotating counterparts (Einsel & Spurzem, 1999). Therefore, the internal structural parameters such as the $r_{\mathrm{h}}$ and $r_{\mathrm{c}}$ change in this initialization step from their original McLuster N-body distribution (see already Fig. 3).
@@
We use a flat IMF between 8.0 and 300.0 $\mathrm{M}_{\odot}$ (Lazar & Bromm, 2022), which gives an initial crossing time of 0.0074 Myr and half-mass relaxation time of 6.869 Myr. Primordial binaries are paired with uniform mass ratios (0.1<q<1.00.1<q<1.0) [^1] The Kobulnicky et al. (2014) distribution adopts a random pairing for stars with $m\leq 5\penalty 10000\ \mathrm{M}_{\odot}{}$, but due to our IMF setup, all stars have $m>5\penalty 10000\ \mathrm{M}_{\odot}{}$. following previous studies (Kiminki et al., 2012; Sana & Evans, 2011; Sana et al., 2012, 2013; Kobulnicky et al., 2014); their semi-major axes are uniformly distributed between 10 and 100 au and their eccentricities are thermal.
@@
We adopt a primordial binary fraction of $f_{\mathrm{b}}=1\%$. A higher $f_{\mathrm{b}}$ at $N\simeq 10^{5}$ is computationally expensive. The Pop III binary fraction remains highly uncertain and likely depends on formation environment and feedback; hydrodynamic simulations can yield high multiplicities (tens of percent; e.g. Stacy & Bromm 2013), while other physical effects may reduce fragmentation and thus the multiplicity (e.g., magnetic field, Sharda & Menon 2025; UV, Hosokawa et al. 2016). We therefore treat $f_{\mathrm{b}}=1\%$ as a limiting assumption that isolates the role of rotation in a controlled pilot suite and provides a conservative lower bound on contributions from primordial binary evolution channels (Type-I and Type-III; Sect. 4.1). We discuss the expected qualitative impact of higher $f_{\mathrm{b}}$ on IMBH formation in Sect. 5.
The cluster metallicity is $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, following Wang et al. (2022) and matching the lowest currently available metallicity for the fitting formulae of Tanikawa et al. (2020).
@@
We place the clusters on a circular orbit at 13.3 kpc following Kamlah et al. (2022b); this reproduces the mass-loss evolution of the eccentric NGC3201 orbit (8.60–29.25 kpc, $e=0.55$, Gaia Collaboration et al., 2018), assuming a point-mass MW of $1.78\times 10^{11}$ $\mathrm{M}_{\odot}$ and $v_{c}=240.0$ $\mathrm{km\penalty 10000\ s}^{-1}$ (Gaia Collaboration et al., 2018; Bobylev & Bajkova, 2020), yielding an initial tidal radius of 264.50 pc. This setup follows Kamlah et al. (2022a, b) to facilitate direct comparison. Our models adopt a static, Milky-Way–like tidal field as a controlled test; we emphasize that realistic high-z environments can produce time-dependent tidal tensors and repeated shocks – effects which cosmological studies show can materially increase mass loss and modify angular-momentum evolution (e.g., Renaud & Gieles, 2015). A complementary study embedding our clusters in cosmological tidal histories is warranted.
@@
We stop the simulations at 500 Myr because (i) the gravothermal–gravogyro catastrophe has largely subsided and the cluster has undergone core-collapse; (ii) most ZAMS stars, particularly the most massive, have reached their final stages; and (iii) environmental effects (e.g. cluster mergers) make long-term isolated evolution of Pop III clusters unlikely (see, e.g., Arca-Sedda & Gualandris, 2018). Thus runs much longer than $\sim$10 Gyr (e.g. Wang et al. 2022) are not always required.
@@
Particles are removed once they reach twice the current tidal radius from the density centre; they are then counted as “escapers”. The tidal radius is recomputed from the current cluster mass.
Apart from the common setup above, we run four initial rotation parameters $\omega_{0}$ with and without GRK, for eight simulations in total (Table 2).
@@ 4.1 Stars and compact objects
Figure 1: Initial-final-mass relation (IFMR) for the NoK0.0 run: black-hole mass $m_{\mathrm{BH}}$ versus progenitor ZAMS mass $m_{\mathrm{ZAMS}}$. The panels correspond to progenitors exploding as lCCSNe, PPISNe, PISNe, and hCCSNe. (IM)BH formation channels are shown in red.
Fig. 5 tracks counts of different stellar types. Counts are similar during the first Myr and then diverge as stellar evolution and core collapse proceed [^2] Core collapse time is $\sim$50-80 Myr depending on $\omega_{0}$; see Sect. 4.3.1.. Larger $\omega_{0}$ speeds stellar loss, most clearly in (No)K1.8, while same-$\omega_{0}$ K/NoK pairs remain nearly identical. The trend reflects the gravothermal-gravogyro catastrophe combined with tidal mass loss. By $\sim$20 Myr all MS stars have evolved and none remain. Near 20 Myr all models produce ¿1 000 NSs ($n_{\mathrm{NS}}$), which later decline mainly because of escapers (Sect. 4.5). BHs appear earlier ($\sim$6 Myr) and $n_{\mathrm{BH}}$ peaks at $\sim$50 000 before 10 Myr.
Fig. 1 and 7 show the initial-final-mass relation (IFMR) for BHs retained in the cluster at 500 Myr, with $m_{\mathrm{BH}}$ ($\mathrm{M}_{\odot}$) versus progenitor ZAMS mass $m_{\mathrm{ZAMS}}$ ($\mathrm{M}_{\odot}$). Progenitors produce BHs via lCCSNe (Fryer et al., 2012), PPISNe (Belczynski et al., 2016), or hCCSNe (Fryer et al., 2012); PISNe leave no remnant (Belczynski et al., 2016) [^3] Also see Appendix C.3 and Kamlah et al. (2022a).. Star-star and star-BH mergers produce IMBHs inside the gap; IMBHs are found above and below it.
We distinguish remnant-mass regimes of (IM)BHs produced by single versus binary evolution. Fig. 1 shows these regimes for the NoK0.0 simulation.
  • Type-I: BHs from primordial binaries where H-envelopes of one or both members are fully stripped (by tides, RLOF, or CEE; see Appendix C.3), leaving naked HeMS cores that collapse to BHs. No mergers occur in this channel (c.f. Type-III); survivors may form a BH-BH binary depending on fallback and natal kicks.
@@
  • Type-III: BHs from primordial-binary mergers (CHeB+MS or CHeB+CHeB during RLOF); the merger product becomes a ShHeB star that later collapses to a BH.
@@
  • Type-V: IMBHs from the PPISNe channel (see Fig. 7; e.g. K0.0).
@@
We identify the formation channels present in the simulations using the Pop III fitting formulae of Tanikawa et al. (2020) and our initial conditions. Full N-body dynamics (instead of population synthesis) is crucial: many (IM)BHs result from interactions in dynamically formed binaries.
Within $\leq$20 Myr and during the pre-core collapse [^4] Core collapse is at 50–80 Myr; see Fig. 8, the Pop III clusters’ stellar population is radically altered. Compared to similar Pop I clusters (Kamlah et al., 2022b), this evolution is more extreme due to the extreme initial conditions. All Pop III cluster models produce (IM)BHs below, within, and above the pair-instability gap.
@@ 4.2 Compact binary fractions and general properties
Figure 2: Binary distributions at 500 Myr. From top to bottom: $e^{2}$, cumulative semi-major axis $a$ (au), cumulative binary potential energy $E_{\mathrm{pot}}$ (N-body units), cumulative distance to the density centre $r_{\mathrm{dens}}$ (pc), and mass ratio $q\equiv m_{2}/m_{1}$. In the $r_{\mathrm{dens}}$ panel a blue line shows the half-mass radius $r_{\mathrm{h}}$ of the NoK0.0 model at 500 Myr ($\sim$9.77 pc).
@@
Fig. 2 summarizes binaries within the tidal radius at 500 Myr. The $e^{2}$ distributions are bottom-heavy and similar across simulations. The cumulative $a$ distributions diverge around $a\approx 50$ au: most binaries are within $a\approx 50$ au in NoK1.8 but not in NoK0.6 or NoK1.2, indicating stronger disruption of wide binaries in the fastest-rotating model mostly before core-collapse (see Fig. 6).
The $E_{\mathrm{pot}}$ distributions show fewer hard binaries in NoK1.8. This can be explained by (i) prior collisions/coalescences (Sect. 4.4; Fig. 6; Table 3; Fig. 9) and/or (ii) preferential escape of hard binaries in rapidly rotating runs (Fig. 10). This effect ejects many would-be merging binaries early, lowering in-situ merger rates of stars and BHs.
@@
In the (No)K0.6 runs GRK reduces binaries by $\sim$50 (mainly at a=4050a=40-50 au, $r_{\mathrm{dens}}=10$ pc, $E_{\mathrm{pot}}=-1000$ N-body units), while in (No)K1.2 it increases binaries by $\sim$30 (mainly at a=5060a=50-60 au, $r_{\mathrm{dens}}=30$ pc). Effects of GRK in the non-rotating and fastest-rotating models are less visible. Overall, our results do not reveal a clear or systematic effect of GRK on the binary distribution within the simulated clusters.
@@ 4.3.1 Structural parameter evolution
Figure 3: Plot showing the total cluster mass $M_{\mathrm{cl}}{}$, the tidal radius $r_{\mathrm{tid}}$, the half mass radius $r_{\mathrm{h}}$, the mass of the core $m_{\mathrm{c}}{}$ and the radius of the core $r_{\mathrm{c}}$ in the four panels for all eight simulations with and without GR recoil kicks for different $\omega_{0}$. Time is shown on a logarithmic scale to highlight the cluster’s rapid early evolution. The models with K are plotted as solid curves and the models without (NoK models) are plotted as dash-dotted curves.
@@
We run each of the four initial models ($\omega_{0}=0.0,0.6,1.2,1.8)$ using Nbody6++GPU with GRK switched on (K models) and without (NoK models). Models with the same $\omega_{0}$ evolve similarly whether or not GRK are enabled, so the dominant control parameter of the global evolution is the initial bulk rotation. Fig. 3 and Fig. 8 show a four-phase evolution:
@@
  • Phase I: before the stellar evolution mass loss begins to dominate the star cluster evolution, namely, 020-2 Myr. $M_{\mathrm{cl}}$ and $r_{\mathrm{tid}}$ stay constant, because mass loss through escapers or stellar evolution is tiny. Depending on $\omega_{0}$, the remaining structural parameter in Fig. 3 show many more variations. Kamlah et al. (2022b) used the same rotating King model setups for Pop I clusters, and showed that angular-momentum transport is very rapid and usually occurs before core collapse; the larger $\omega_{0}$, the faster this phenomenon happens. In the Pop III clusters in this work, we also observe this fast angular momentum transport, for example in the panel showing the half-mass radius of Fig. 3. Within the first Myr, the larger $\omega_{0}$, the more pronounced and earlier is the temporary increase of $r_{\mathrm{h}}$ and decrease shortly after. This evolution is even more apparent in Fig. 8, which shows the evolution of the Lagrangian radii $r_{\mathrm{Lagr}}$ within spheres that contain 1%, 5%, 10%, 30%, 50%, and 90% of the total cluster mass at the current simulation time-step. For example, in the (No)K1.8 models, the degree of initial bulk rotation has a pronounced effect: the outermost $r_{\mathrm{Lagr}}$ expands substantially, while the other Lagrangian radii generally contract up to 2 Myr. All $r_{\mathrm{Lagr}}$ show a small and synchronous bump (i.e., a brief increase followed by a decrease) at around 0.1 Myr. The (No)K1.2 models show the same effect, but weaker and slightly delayed relative to the (No)K1.8 models. Such a phenomenon is not shown in (No)K0.0 and (No)K0.6 models, where $r_{\mathrm{h}}$ stays approximately constant until 2 Myr. This early evolution is also mirrored by $r_{\mathrm{c}}$ and $m_{\mathrm{c}}$, where both the (No)K1.2 and the (No)K1.8 exhibit fast increases with subsequent decreases in the respective models. Again, this evolution is more pronounced and happens earlier in the (No)K1.8 with respect to the (No)K1.2 models. In summary, from 0 to $2$ Myr, the evolution is governed only by the initial bulk rotation, because tidal and stellar-evolution mass losses are negligible.
[…]
@@
  • Phase III: from the end of phase II ($\sim$3 Myr) to the time of gravothermal core-collapse. As shown in Fig. 8, the ending time of this phase varies between models, which is $\sim$80 Myr for (No)K0.0, $\sim$70 Myr for (No)K0.6, $\sim$60 Myr for (No)K1.2, and $\sim$50 Myr for (No)K1.8. Therefore, the larger the degree of initial bulk rotation in our Pop III clusters, the higher the central density, and the shorter the gravothermal collapse timescale, as seen in previous studies (e.g., Akiyama & Sugimoto, 1989; Einsel & Spurzem, 1999; Kim et al., 2002, 2004, 2008; Fiestas et al., 2012; Kamlah et al., 2022b). Larger $\omega_{0}$ leads to higher $M_{\mathrm{av}}$ in the 5% and 10% spheres at core collapse. Thus, stronger initial rotation accelerates mass segregation. In (No)K1.8 models, the cluster expansion ejects many stars [^5] See the next paper in this series., producing a decrease of the 90% $r_{\mathrm{Lagr}}$.

### 13 Apr 2026 v2

@@ 2 Methods
We used Nbody6++GPU (Spurzem, 1999; Nitadori & Aarseth, 2012; Wang et al., 2015) for the dynamical evolution of the Pop III clusters. The initial rotating conditions followed Kamlah et al. (2022b) and were generated with McLuster (Kuepper et al., 2011; Leveque et al., 2022b) and Fopax (Einsel & Spurzem, 1999); details are given in Appendix C.2.
Population III stellar evolution was treated with the low-metallicity recipes of Tanikawa et al. (2020) at $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, without radiation-, pulsation- (e.g. Nakauchi et al., 2020), or rotation-driven (e.g. Liu et al., 2021b) wind mass loss; further details are in Appendix C.3. Fitting formulae for relativistic merger recoil kick velocity and remnant mass and spin followed Campanelli et al. (2007); Lousto et al. (2012); Jiménez-Forteza et al. (2017), as summarized in Appendix C.4. We refer to BHs with masses above 100 $\mathrm{M}_{\odot}$ as IMBHs.
@@ 3 Initial conditions
The mass function of Pop III star clusters is highly uncertain. Formation of Pop III star clusters has been simulated in cosmological mini-halos (cf. e.g. Stacy et al., 2016), and corresponding star-cluster simulations have covered masses from $10^{3}$ to $10^{5}$ $\mathrm{M}_{\odot}$ (Liu et al., 2025; Wu et al., 2025). However, Kimm et al. (2016) propose the formation of more massive Pop III clusters without dark matter directly from cooling gas clouds, with masses of up to $5\cdot 10^{5}\mathrm{M}_{\odot}{}$, and Liu et al. (2021a) find cluster masses reaching this value. Stacy et al. (2016) also discuss a possible Pop III cluster with a mass as high as $10^{7}\mathrm{M}_{\odot}{}$. In our simulations we explore the very high-mass end (over $10^{5}$ objects, with 1% binaries and a top-heavy IMF; see Table 1) for two reasons: first, to extend previously published simulations to higher masses; and second, to probe very high initial densities that may promote rapid formation of massive stars and massive BHs (Vergara et al., 2025).
@@
We started from an unsegregated, unfractal King model (King, 1966; Goodwin & Whitworth, 2004) with $W_{0}=6.0$, initially in virial equilibrium. The half-mass radius was set to $r_{\mathrm{h}}$=1 pc. As is outlined in Appendix C.2, the initial model from McLuster was then redistributed with a rotating King model, which is more compact than its non-rotating counterparts (Einsel & Spurzem, 1999). Therefore, the internal structural parameters such as the $r_{\mathrm{h}}$ and $r_{\mathrm{c}}$ changed in this initialization step from their original McLuster N-body distribution (see Fig. 3).
@@
We used a flat IMF of between 8.0 and 300.0 $\mathrm{M}_{\odot}$ (Lazar & Bromm, 2022), which gave an initial crossing time of 0.0074 Myr and half-mass relaxation time of 6.869 Myr. Primordial binaries were paired with uniform mass ratios (0.1<q<1.00.1<q<1.0) [^1] The Kobulnicky et al. (2014) distribution adopts a random pairing for stars with $m\leq 5\penalty 10000\ \mathrm{M}_{\odot}{}$, but due to our IMF set-up, all stars have $m>5\penalty 10000\ \mathrm{M}_{\odot}{}$. following previous studies (Kiminki et al., 2012; Sana & Evans, 2011; Sana et al., 2012, 2013; Kobulnicky et al., 2014); their semi-major axes were uniformly distributed between 10 and 100 au and their eccentricities were thermal. The cluster metallicity was $Z/\mathrm{Z}_{\odot}{}=10^{-8}$, following Wang et al. (2022) and matching the lowest currently available metallicity for the fitting formulae of Tanikawa et al. (2020).
@@
We adopted a primordial binary fraction of $f_{\mathrm{b}}=1\%$. A higher $f_{\mathrm{b}}$ at $N\simeq 10^{5}$ is computationally expensive. The Pop III binary fraction remains highly uncertain and likely depends on formation environment and feedback; hydrodynamic simulations can yield high multiplicities (tens of percent; e.g. Stacy & Bromm 2013), while other physical effects may reduce fragmentation, and thus the multiplicity (e.g. magnetic field; Sharda & Menon 2025; UV, Hosokawa et al. 2016). We therefore treated $f_{\mathrm{b}}=1\%$ as a limiting assumption that isolates the role of rotation in a controlled pilot suite and provides a conservative lower bound on contributions from primordial binary evolution channels (Type-I and Type-III; Sect. 4.1). We discuss the expected impact of higher $f_{\mathrm{b}}$ on IMBH formation in Sect. 5.
@@
We placed the clusters on a circular orbit at 13.3 kpc following Kamlah et al. (2022b); this reproduced the mass-loss evolution of the eccentric NGC3201 orbit (8.60–29.25 kpc, $e=0.55$, Gaia Collaboration et al., 2018), assuming a point-mass MW of $1.78\times 10^{11}$ $\mathrm{M}_{\odot}$ and $v_{c}=240.0$ $\mathrm{km\penalty 10000\ s}^{-1}$ (Gaia Collaboration et al., 2018; Bobylev & Bajkova, 2020), yielding an initial tidal radius of 264.50 pc. This set-up followed Kamlah et al. (2022a, b) to facilitate a direct comparison. Our models adopted a static, Milky-Way–like tidal field as a controlled test; we emphasize that realistic high-z environments can produce time-dependent tidal tensors and repeated shocks – effects that cosmological studies show can materially increase mass loss and modify angular-momentum evolution (e.g. Renaud & Gieles, 2015). A complementary study embedding our clusters in cosmological tidal histories is warranted.
@@
We stopped the simulations at 500 Myr because (i) the gravothermal–gravogyro catastrophe has largely subsided and the cluster has undergone core-collapse; (ii) most ZAMS stars, particularly the most massive ones, have reached their final stages; and (iii) environmental effects (e.g. cluster mergers) make long-term isolated evolution of Pop III clusters unlikely (see e.g. Arca-Sedda & Gualandris, 2018). Thus runs much longer than $\sim$10 Gyr (e.g. Wang et al. 2022) are not always required.
@@
Particles were removed once they reached twice the current tidal radius from the density centre; they were then counted as ‘escapers’. The tidal radius was recomputed from the current cluster mass. Apart from the common set-up above, we ran four initial rotation parameters, $\omega_{0}$, with and without GRKs, for eight simulations in total (Table 2).
@@ 4.1 Stars and compact objects
Figure 1: Initial-final-mass relation (IFMR) for the NoK0.0 run: BH mass, $m_{\mathrm{BH}}$, versus progenitor ZAMS mass, $m_{\mathrm{ZAMS}}$. The panels correspond to progenitors exploding as lCCSNe, PPISNe, PISNe, and hCCSNe. (IM)BH formation channels are shown in red.
Figure 5 tracks counts of different stellar types. Counts are similar during the first million years and then diverge as stellar evolution and core collapse proceed [^2] Core collapse time is $\sim$50-80 Myr depending on $\omega_{0}$; see Sect. 4.3.1.. Larger $\omega_{0}$ speeds stellar loss, most clearly in (No)K1.8, while same-$\omega_{0}$ K/NoK pairs remain nearly identical. The trend reflects the gravothermal-gravogyro catastrophe combined with tidal mass loss. By $\sim$20 Myr all MS stars have evolved and none remain. Near 20 Myr all models produce ¿1 000 NSs ($n_{\mathrm{NS}}$), which later decline mainly because of escapers (Sect. 4.5). The BHs appear earlier ($\sim$6 Myr) and $n_{\mathrm{BH}}$ peaks at $\sim$50 000 before 10 Myr.
We divided the ZAMS particles into four classes: low-mass core-collapse supernovae (lCCSNe): $8.0\leq m_{\mathrm{ZAMS}}<85.0\penalty 10000\ \mathrm{M}_{\odot}$, pulsational pair-instability supernovae (PPISNe): $85.0\leq m_{\mathrm{ZAMS}}<120.0\penalty 10000\ \mathrm{M}_{\odot}$, pair-instability supernovae (PISNe): $120.0\leq m_{\mathrm{ZAMS}}<265.0\penalty 10000\ \mathrm{M}_{\odot}$, and high-mass core-collapse supernovae (hCCSNe): $265.0\leq m_{\mathrm{ZAMS}}<300.0\penalty 10000\ \mathrm{M}_{\odot}$, where $m_{\mathrm{ZAMS}}$ is the ZAMS stellar mass of a single star [^3] A primordial binary may have members of two different mass groups.. Figures 1 and 7 show the initial-final-mass relation (IFMR) for BHs retained in the cluster at 500 Myr, with $m_{\mathrm{BH}}$ ($\mathrm{M}_{\odot}$) versus the progenitor ZAMS mass, $m_{\mathrm{ZAMS}}$ ($\mathrm{M}_{\odot}$). Progenitors produce BHs via lCCSNe (Fryer et al., 2012), PPISNe (Belczynski et al., 2016), or hCCSNe (Fryer et al., 2012); PISNe leave no remnant (Belczynski et al., 2016) [^4] See also Appendix C.3 and Kamlah et al. (2022a).. Star-star and star-BH mergers produce IMBHs inside the gap; IMBHs are found above and below it.
We distinguish remnant-mass regimes of (IM)BHs produced by single versus binary evolution. Figure 1 shows these regimes for the NoK0.0 simulation [^5] Stellar type acronyms follow Hurley et al. (2000); also see Sect. C.3. .
  • Type-I: The BHs from primordial binaries where H-envelopes of one or both members are fully stripped (by tides, RLOF, or CEE; see Appendix C.3), leaving naked HeMS cores that collapse into BHs. No mergers occur in this channel (c.f. Type-III); survivors may form a BH-BH binary depending on fallback and natal kicks.
@@
  • Type-III: The BHs from primordial-binary mergers (CHeB+MS or CHeB+CHeB during RLOF); the merger product becomes a ShHeB star that later collapses to a BH.
@@
  • Type-V: The IMBHs from the PPISN channel (see Fig. 7; e.g. K0.0).
@@
We identified the formation channels present in the simulations using the Pop III fitting formulae of Tanikawa et al. (2020) and our initial conditions. Full N-body dynamics (instead of population synthesis) is crucial: many (IM)BHs result from interactions in dynamically formed binaries.
Within $\leq$20 Myr and during the pre-core collapse [^6] Core collapse is at 50–80 Myr; see Fig. 8, the Pop III clusters’ stellar population is radically altered. Compared to similar Pop II clusters (Kamlah et al., 2022b), this evolution is more extreme due to the extreme initial conditions. All Pop III cluster models produce (IM)BHs below, within, and above the pair-instability gap.
@@ 4.2 Compact binary fractions and general properties
Figure 2: Binary distributions at 500 Myr. From top to bottom: $e^{2}$; cumulative semi-major axis, $a$ (au); cumulative binary potential energy, $E_{\mathrm{pot}}$ (N-body units); cumulative distance to the density centre, $r_{\mathrm{dens}}$ (parsecs); and mass ratio $q\equiv m_{2}/m_{1}$. In the $r_{\mathrm{dens}}$ panel a blue line shows the half-mass radius $r_{\mathrm{h}}$ of the NoK0.0 model at 500 Myr ($\sim$9.77 pc).
@@
Figure 2 summarizes binaries within the tidal radius at 500 Myr. The $e^{2}$ distributions are bottom-heavy and similar across simulations. The cumulative $a$ distributions diverge around $a\approx 50$ au: most binaries are within $a\approx 50$ au in NoK1.8 but not in NoK0.6 or NoK1.2, indicating stronger disruption of wide binaries in the fastest-rotating model mostly before core-collapse (see Fig. 6).
The $E_{\mathrm{pot}}$ distributions show fewer hard binaries in NoK1.8. This can be explained by (i) prior collisions/coalescences (Sect. 4.4, Figs. 6 and 9, Table 3) and/or (ii) preferential escape of hard binaries in rapidly rotating runs (Fig. 10). This effect ejects many would-be merging binaries early, lowering in situ merger rates of stars and BHs.
@@
In the (No)K0.6 runs GRK reduces binaries by $\sim$50 (mainly at a=4050a=40-50 au, $r_{\mathrm{dens}}=10$ pc, $E_{\mathrm{pot}}=-1000$ N-body units), while in (No)K1.2 it increases binaries by $\sim$30 (mainly at a=5060a=50-60 au, $r_{\mathrm{dens}}=30$ pc). The effects of GRKs in the non-rotating and fastest-rotating models are less visible. Overall, our results do not reveal a clear or systematic effect of GRKs on the binary distribution within the simulated clusters.
@@ 4.3.1 Structural parameter evolution
Figure 3: Total cluster mass, $M_{\mathrm{cl}}{}$, tidal radius, $r_{\mathrm{tid}}$, half mass radius, $r_{\mathrm{h}}$, mass of the core, $m_{\mathrm{c}}{}$, and radius of the core, $r_{\mathrm{c}}$, in the four panels for all eight simulations with and without GR recoil kicks for different $\omega_{0}$. Time is shown on a logarithmic scale to highlight the cluster’s rapid early evolution. The models with K are plotted as solid curves and the models without (NoK models) are plotted as dash-dotted curves.
@@
We ran each of the four initial models ($\omega_{0}=0.0,0.6,1.2,1.8)$ using Nbody6++GPU with GRKs switched on (K models) and without (NoK models). Models with the same $\omega_{0}$ evolve similarly whether or not GRKs are enabled, so the dominant control parameter of the global evolution is the initial bulk rotation. Figures 3 and 8 show a four-phase evolution:
@@
  • Phase I: Before the stellar evolution mass loss begins to dominate the star cluster evolution, namely, 020-2 Myr. $M_{\mathrm{cl}}$ and $r_{\mathrm{tid}}$ stay constant, because mass loss through escapers or stellar evolution is tiny. Depending on $\omega_{0}$, the other structural parameters in Fig. 3 show many more variations. Kamlah et al. (2022b) used the same rotating King model set-ups for Pop II clusters, and showed that angular-momentum transport is very rapid and usually occurs before core collapse; the larger $\omega_{0}$, the faster this phenomenon happens. In the Pop III clusters in this work, we also observe this fast angular momentum transport, for example in the panel showing the half-mass radius of Fig. 3. Within the first million years, the larger $\omega_{0}$, the more pronounced and earlier is the temporary increase in $r_{\mathrm{h}}$ and decrease shortly after. This evolution is even more apparent in Fig. 8, which shows the evolution of the Lagrangian radii $r_{\mathrm{Lagr}}$ within spheres that contain 1%, 5%, 10%, 30%, 50%, and 90% of the total cluster mass at the current simulation time step. For example, in the (No)K1.8 models , the degree of initial bulk rotation has a pronounced effect: the outermost $r_{\mathrm{Lagr}}$ expands substantially, while the other Lagrangian radii generally contract up to 2 Myr. All $r_{\mathrm{Lagr}}$ show a small and synchronous bump (i.e. a brief increase followed by a decrease) at around 0.1 Myr. The (No)K1.2 models show the same effect, but weaker and slightly delayed relative to the (No)K1.8 models. Such a phenomenon is not shown in (No)K0.0 and (No)K0.6 models, where $r_{\mathrm{h}}$ stays roughly constant until 2 Myr. This early evolution is also mirrored by $r_{\mathrm{c}}$ and $m_{\mathrm{c}}$, where both the (No)K1.2 and the (No)K1.8 exhibit fast increases with subsequent decreases in the respective models. Again, this evolution is more pronounced and happens earlier in the (No)K1.8 with respect to the (No)K1.2 models. In summary, from 0 to $2$ Myr, the evolution is governed only by the initial bulk rotation; tidal and stellar-evolution mass losses are negligible.
  • Phase II: Rapid stellar evolution mass loss around 232-3 My r, as shown in 1st panel of Fig. 3. This is mainly due to compact object formation, particularly PISN events, which produce no BHs and lose all the mass. Besides stellar evolution, mass loss is also caused by (a) escaping stars, through compact object natal kicks or GRKs, or (b) dynamical encounters, including strong encounters or by cumulative weak encounters. The cluster kinematics also change here owing to fallback-dependent natal kicks, which are generally larger for lower-mass BHs (see Kamlah et al., 2022a). As a result of the total mass drop, $r_{\mathrm{tid}}$ drops, and $r_{\mathrm{h}}$ with $r_{\mathrm{c}}$ rises sharply in all models. Interestingly, the maxima of $r_{\mathrm{c}}$ lie in the (No)K1.2 model, while the cores of the (No)K1.8 and (No)K0.0 models are similarly compact. We compare this with the average mass in the core $M_{\mathrm{av}}$ within different Lagrangian radii $r_{\mathrm{Lagr}}$, shown in Fig. 8. Signified by $M_{\mathrm{av}}$, we see that here all models show a similar segregation of masses into different spheres independent of initial bulk rotation; for example, the minimum $M_{\mathrm{av}}$ in the 10% $r_{\mathrm{Lagr}}$ has similar values occurring at similar times, although we note that this quantity is subject to large statistical fluctuations.
@@
  • Phase III: From the end of phase II ($\sim$3 Myr) to the time of gravothermal core-collapse. As shown in Fig. 8, the ending time of this phase varies between models, which is $\sim$80 Myr for (No)K0.0, $\sim$70 Myr for (No)K0.6, $\sim$60 Myr for (No)K1.2, and $\sim$50 Myr for (No)K1.8. Therefore, the larger the degree of initial bulk rotation in our Pop III clusters, the higher the central density, and the shorter the gravothermal collapse timescale, as seen in previous studies (e.g. Akiyama & Sugimoto, 1989; Einsel & Spurzem, 1999; Kim et al., 2002, 2004, 2008; Fiestas et al., 2012; Kamlah et al., 2022b). Larger $\omega_{0}$ leads to higher $M_{\mathrm{av}}$ in the 5% and 10% spheres at core collapse. Thus, stronger initial rotation accelerates mass segregation. In (No)K1.8 models, the cluster expansion ejects many stars [^7] See the next paper in this series., producing a decrease in the 90% $r_{\mathrm{Lagr}}$.